# Language Model Alignment in Multilingual Trolley Problems

### v4

@@ 7 Conclusion
Our study provides a comprehensive analysis of the moral preferences exhibited by LLMs across a wide range of languages, via the prism of trolley problems. We propose a multilingual large-scale dataset, MultiTP, with systematic variation of the vignette design. Our dataset is also equipped with an unprecedentedly large set of human responses of 40 million responses from over 200 countries. Our experiments assess the moral alignment of 19 LLMs across 100+ languages, and find that most LLMs do not demonstrate strong alignment with human preferences on trolley problem questions, but in the meantime there is no significant inequality across languages, as high-resource languages have similar alignment scores to low-resource languages. Our study paves the way for future pluralistic alignment research grounded in psychology and more languages.
Reproducibility Statement
We have put efforts to ensure the reproducibility of this work. Our MultiTP data is uploaded to the submission system and we have described our data construction details in Section 3 as well as the overall statistics. In the meantime, we have also uploaded our code to the submission system. Our Section 4 in the main paper describes our model setup and the equations used for the evaluation metrics. We provide an extensive appendix to supplement all the rest of the details.
Ethics Statement
In studying the six moral dimensions of species, gender, fitness, status, age, and number, we acknowledge that our work is not intended to promote the integration of LLMs into systems requiring critical real-world judgments, such as autonomous driving or other high-stakes scenarios. Instead, our research aims to use these moral dilemmas as a tool of theoretical exploration for better understanding LLM alignment with human values. For instance, while sparing humans over animals is often seen as an intuitive moral choice in many cultures, moral values can vary significantly across underrepresented groups, complicating the assumption of universally “correct” answers. Thus, it is essential for LLM practitioners to remain vigilant in ensuring that embedded assumptions reflect a wide range of cultural and linguistic contexts.
We also recognize the importance of broadening the types of moral dilemmas used to assess LLM alignment. While our research focuses on the trolley problem, a well-known ethical dilemma, we are not suggesting this singular scenario captures the complexity of moral reasoning. Expanding the scope to include dilemmas such as organ transplants or firefighter decisions can offer richer insights into how LLMs process ethical judgments. By evaluating LLMs on a diverse set of moral challenges, we can gain a more complete picture of their alignment capabilities, rather than over-relying on one specific framework.
Additionally, we understand the limitations of using binary moral dilemmas, such as trolley problems, to evaluate LLM decision-making. Reducing moral judgments to binary choices risks oversimplifying the complex, context-dependent nature of human morality (Schein, 2020). While these scenarios provide valuable insights into LLM behavior, they should not be seen as definitive measures of moral competence. Our research emphasizes that LLMs, as currently developed, are not suitable for real-world moral decision-making tasks. Rather, this work serves as a means to explore how LLMs can better align with human moral frameworks in a controlled, research-oriented environment. We encourage further research to ensure that AI systems are fair, transparent, and accountable as they develop, but caution against their use in high-stakes moral applications without substantial ethical oversight.
@@ B.2 Country-to-Language Mapping
To allow for such analysis, we collect each country and their main languages from Wikipedia population statistics. We curate the list manually by going through the Wikipedia page of the languages of each country, such as this one for Belgium: https://en.wikipedia.org/wiki/Languages_of_Belgium. To account for multilingual speakers, we use the first-language only speaker information, i.e., the number of speakers who speaks the language as their first language.
If future researchers want to calculate the country-specific misalignment scores, we suggest a weighted average by each language-specific MIS score and the number of speakers of that language in that country.
@@ C.1 Model Setup
Across all our LLM evaluation experiments, to ensure reproduceability, we set the text generation temperature to zero for decoding.
We first introduce all the exact model identifiers for the open-weights models in Table 5. Then, we provide the model ID for close-weights models using the OpenAI API in Table 5. The total estimate API cost for the experiment is 600 USD.
Table 4: Detail model identifier and VRAM requirements when running open-weights models.
@@
Model Size Cost per Language Identifier
@@ C.2 Prompt Paraphrases
Appendix D Additional Results
D.1 LLM Clustering Results
@@
Figure 5: Distribution of preferences by feature across languages for GPT-3
Table 8: Language in each cluster for GPT-3
Cluster Language
@@ D Welsh, German, Greek, Hebrew, Japanese, Korean, Luxembourgish, Swahili, Filipino, Turkish, Ukrainian
Figure 6: Distribution of preferences by feature across languages for GPT-4
Table 9: Language in each cluster for GPT-4
cluster language
@@ D Cebuano, Spanish, Latin, Tajik
D.2 Correlation of the Misalignment Score and the Number of Speakers of Each Language
@@
D.3 Country-Specific Alignment
@@
Figure 7: Moral misalignment world map. The shade of the color corresponds to misalignment score of Llama 3.1 70B, where a darker color indicates larger misalignment. We aggregate the country-specific alignment or according to the approximation procedures introduced in Section B.2.

### v6

@@ 8 Conclusion
Our study provides a comprehensive analysis of the moral preferences exhibited by LLMs across a wide range of languages, via the prism of trolley problems. We propose a multilingual large-scale dataset, MultiTP, with systematic variations of the vignette design. Our dataset is also equipped with an unprecedentedly large set of human responses of 40 million responses from over 200 countries. Our experiments assess the moral alignment of 19 LLMs across 100+ languages, and find that most LLMs do not demonstrate strong alignment with human preferences on trolley problem questions, but in the meantime, there is no significant inequality across languages, as high-resource languages have similar alignment scores to low-resource languages. Our study paves the way for future pluralistic alignment research grounded in psychology and more languages.
Acknowledgments
This work was partially supported by a grant to SL from the Templeton World Charity Foundation, TWCF-2023-32585. This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645. The usage of OpenAI credits is largely supported by the Tübingen AI Center. MKW was supported by the Cooperative AI foundation and Foresight Institute.
@@ B.2 Country-to-Language Mapping
To allow for such analysis, we collect each country and its main languages from Wikipedia population statistics. We curate the list manually by going through the Wikipedia page of the languages of each country, such as this one for Belgium: https://en.wikipedia.org/wiki/Languages_of_Belgium. To account for multilingual speakers, we use the first-language-only speaker information, i.e., the number of speakers who speak the language as their first language.
If future researchers want to calculate the country-specific misalignment scores, we tentatively suggest a weighted average by each language-specific MIS score and the number of speakers of that language in that country.
@@ C.1 Model Setup
To ensure reproducibility, we set the text generation temperature to zero for greedy decoding across all our LLM evaluation experiments.
Table 5 includes all the exact model identifiers for the open-weights models, and also the model ID for close-weights models using the OpenAI API in Table 5. The total estimated API cost for the experiment is around 600 USD.
Table 4: Detail model identifiers and VRAM requirements for the open-weights models.
@@
Model Version Cost per Language Identifier
@@ C.2 Prompt Paraphrases
Appendix D Translation Quality Check
D.1 Human Evaluation of Translation Quality
To evaluate translation quality, we conducted a human evaluation task on MTurk, where annotators rated how accurately translations from English to another language conveyed the original meaning. Ratings were provided on a 5-point scale ranging from "None" (no meaning conveyed) to "All" (full meaning conveyed). The task was divided into two main phases:
First, annotators were presented with a language selection page. On this page, they were prompted to select the language in which they were most proficient from a list of 107 languages (excluding English). Annotators were explicitly reminded not to proceed if they were unfamiliar with any of the listed languages. This step ensured that the ratings were provided only by individuals fluent in the target languages.
Next, annotators were directed to the translation rating page. This page contained 25 translation pairs, each comprising an English sentence or word and its corresponding translation. The instructions for this phase were adapted from prior research (Lavie, 2011; Goto et al., 2014), as shown below:
Please rate how accurately the translation conveys the meaning of the source text:Source English Sentence/Text: {source}Translation: {translation}- None: No meaning of the original text is conveyed.- Little: Only a small portion of the meaning is conveyed.- Much: A substantial amount of the meaning is conveyed.- Most: Most of the meaning is conveyed accurately.- All: The full meaning is conveyed accurately.
After completing the rating task, annotators submitted their responses, which were then recorded for analysis. We kept the Amazon Mechanical Turk task open for a span of three days, and collected a total of 169 annotation responses covering 44 languages. The total cost was approximately \$10. The languages evaluated included af, am, ar, be, bg, bn, bs, cs, da, de, en, es, fi, fr, ga, gu, hi, hmn, hu, hy, it, ja, ko, la, mg, ml, ms, nl, pl, pt, ro, ru, sq, st, sw, ta, te, tl, tr, ur, vi, yo, zh-cn, and zh-tw. The number of annotations received per language ranged from one response (e.g., am, be, bs, da, fi, gu, hmn, ms, nl, pl, st) to as many as 16 responses (es).
The results of the evaluation demonstrate strong translation quality. Specifically, 88.6% of the evaluated languages achieved an average score of 3.0 or higher out of 5 (i.e., conveying a substantial portion of the original meaning). Importantly, no languages received a mean score below 2.0, and the lowest-performing group only has four languages (am, pl, sq, la), whose mean scores are between 2.0 and 2.5. The standard deviation of scores across languages ranged from 0.5 to 1.3, suggesting relatively consistent ratings among annotators. These findings cover both high-resource and low-resource languages, highlighting the robustness of the evaluation. Overall, the results demonstrate that the translations generally maintain high quality across a diverse set of linguistic contexts. Additionally, we examined the correlation between speaker count and translation score. The Pearson correlation is 0.078, and the Spearman correlation is 0.072, both indicating no meaningful correlation. Given the limited data available, it is reassuring to observe that the relationship remains weak.
Table 7 presents the full annotation results, including details of the 44 evaluated languages, the number of responses received per language, the mean scores, and the standard deviations.
Language Language Code Averaged Scores # Responses
Danish da 5.00 $\pm$0.00 1
Southern Sotho st 5.00 $\pm$0.00 1
Bosnian bs 4.91 $\pm$0.28 1
Turkish tr 4.81 $\pm$0.43 3
Russian ru 4.78 $\pm$0.41 2
Chinese (Traditional) zh-tw 4.76 $\pm$0.87 9
Portuguese pt 4.72 $\pm$0.70 6
German de 4.69 $\pm$0.70 6
Tagalog tl 4.68 $\pm$0.61 4
Swahili sw 4.68 $\pm$1.11 3
Chinese (Simplified) zh-cn 4.63 $\pm$0.98 5
Dutch nl 4.61 $\pm$0.71 1
Korean ko 4.59 $\pm$1.13 3
Bulgarian bg 4.59 $\pm$1.31 2
Czech cs 4.57 $\pm$1.04 2
English en 4.53 $\pm$0.73 7
Malagasy mg 4.52 $\pm$1.12 5
Vietnamese vi 4.52 $\pm$0.50 3
Spanish es 4.49 $\pm$0.96 16
Romanian ro 4.48 $\pm$0.83 2
Belarusian be 4.48 $\pm$0.50 1
Italian it 4.47 $\pm$1.00 5
Japanese ja 4.45 $\pm$1.01 5
Afrikaans af 4.33 $\pm$0.55 2
Malay ms 4.30 $\pm$0.95 1
Telugu te 4.30 $\pm$1.07 3
Yoruba yo 4.29 $\pm$1.25 3
Hungarian hu 4.26 $\pm$1.26 2
Malayalam ml 4.10 $\pm$1.43 5
Tamil ta 4.06 $\pm$1.14 9
Bengali bn 4.03 $\pm$1.14 3
French fr 4.03 $\pm$0.98 3
Hindi hi 3.94 $\pm$1.32 11
Irish ga 3.83 $\pm$1.17 2
Gujarati gu 3.83 $\pm$1.86 1
Urdu ur 3.64 $\pm$1.43 4
Hmong hmn 3.61 $\pm$1.86 1
Arabic ar 3.58 $\pm$1.39 3
Armenian hy 3.28 $\pm$1.27 14
Finnish fi 2.70 $\pm$1.40 1
Latin la 2.50 $\pm$1.31 4
Albanian sq 2.30 $\pm$2.11 2
Polish pl 2.26 $\pm$0.53 1
Amharic am 2.00 $\pm$0.00 1
Table 7: Human evaluation results of translation quality. Annotations were collected for 44 languages, with scores averaged across all translation pairs for each language.
D.2 Automatic Evaluation of Translation Quality
In addition to manual annotation of the translation quality, we also provide the automatic evaluation results by back-translation. Starting with our original prompt in English $x$, we use our paper’s setup to translate and collect the non-English prompt $y$, and based on $y$, we further translate it back to English, thus getting the back-translated English prompt $x^{\prime}$. As a simple method for evaluation, we calculate the cosine similarity of the embeddings between all $x$ and $x^{\prime}$ pairs for all languages, with a histogram plotted in Figure 5.
Figure 5: Distribution of embedding similarity
Appendix E Additional Results
E.1 LLM Clustering Results
@@
Figure 6: Distribution of preferences by feature across languages for GPT-3.
Table 9: Languages in each cluster for GPT-3.
Cluster Languages
@@ D Welsh, German, Greek, Hebrew, Japanese, Korean, Luxembourgish, Swahili, Filipino, Turkish, Ukrainian
Figure 7: Distribution of preferences by feature across languages for GPT-4.
Table 10: Languages in each cluster for GPT-4.
Cluster Languages
@@ D Cebuano, Spanish, Latin, Tajik
E.2 Correlation of the Misalignment Score and the Number of Speakers of Each Language
@@
E.3 Correlation of the Misalignment Score and Language Sensitivity
Table 12 shows the misalignment scores and language sensitivity scores for each model. We observe a moderate positive correlation (Pearson coefficient = 0.43, p-value = 0.07).
Table 12: Misalignment scores and language sensitivity scores for each model.
Model Misalignment Language Sensitivity
Llama 3.1 70B 0.55 18.01
Llama 3 70B 0.56 15.25
Llama 3 8B 0.57 14.90
GPT-3 0.64 15.13
Llama 3.1 8B 0.75 19.92
Qwen 2 7B 0.77 22.23
Mistral 7B 0.80 21.30
GPT-4 0.81 15.83
Llama 2 7B 0.83 19.79
Llama 2 70B 0.91 18.53
Phi-3.5 Mini 0.94 21.27
Gemma 2 2B 0.96 22.88
Phi-3 Medium 1.07 22.79
Phi-3.5 MoE 1.08 14.67
Gemma 2 9B 1.08 24.74
Llama 2 13B 1.10 21.01
Gemma 2 27B 1.17 21.66
Qwen 2 72B 1.20 21.11
GPT-4o Mini 1.45 18.09
E.4 Country-Specific Alignment
@@
Figure 8: Moral misalignment world map of the best model, Llama 3.1 70B. The darker shade of the color corresponds to a larger misalignment score. We aggregate the country-specific alignment or according to the approximation procedures introduced in Section B.2.
E.4.1 Regional Differences across Global East, West, and South
In addition to country-level statistics, we also analyze across three major geographic regions – the Global East, West, and South, using the clustering of Awad et al. (2018; 2020a) – to investigate potential cultural variations in moral decision-making. We show the regional moral preferences of three models, GPT-3, GPT-4, and Llama 3.1 70B, in Figures 9, 10 and 11, respectively.
Figure 9: Distribution of preferences by feature across languages for Llama 3.1 70B for different cultures.
Figure 10: Distribution of preferences by feature across languages for GPT-3 for different cultures.
Figure 11: Distribution of preferences by feature across languages for GPT-4 for different cultures.
E.5 Robustness Analysis: Testing the Option Order Bias
Recent work reports that LLMs often suffer from recency bias (Liu et al., 2024), making them more likely to choose the later option in multiple choice questions.
Since this will have an important effect on our study, we conduct a sanity check to report the consistency rate, i.e., the frequency of LLMs to keep its response if we swap the order the order of the two choices, e.g., mentioning the boy first, and elderly man next, or vice versa in our example in Figure 1. We report the average consistency rates across different models in Table 13, showing that most models are robust against option order changes, with close-to-perfect consistency rates.
Table 13: Consistency rates against the position bias in the option order for the LLMs tested in our study.
Model Consistency Rate
GPT-4o Mini 99.5
Gemma 2 27B 99.2
Gemma 2 9B 99.0
Qwen 2 72B 98.6
GPT-4 98.3
Llama 2 13B 97.7
Mistral 7B 97.3
Llama 2 70B 97.2
Qwen 2 7B 97.1
Phi-3 Medium 96.7
Phi-3.5 MoE 96.6
Llama 2 7B 96.3
Llama 3.1 8B 96.1
Llama 3 70B 95.8
Llama 3.1 70B 95.1
Gemma 2 2B 95.1
Phi-3.5 Mini 94.7
Llama 3 8B 94.1
GPT-3 87.1
E.6 Analysis of Models from the Llama Family
Model Sparing Sparing Sparing Sparing Sparing Sparing
Young Fit Females Higher Status Humans More
Llama 3.1 70B 76.3 70.2 84.6 65.7 87.1 80.4
Llama 3.1 8B 64.2 68.7 72.2 63.7 76.5 78.2
Llama 3 70B 77.3 69.0 91.4 70.9 93.8 83.2
Llama 3 8B 51.6 68.1 84.1 70.3 88.0 73.7
Llama 2 70B 68.7 76.5 83.5 72.2 96.1 79.8
Llama 2 13B 61.7 66.3 86.5 71.5 90.3 76.2
Llama 2 7B 47.7 67.6 84.3 67.7 92.1 78.6
Table 14: Llama family average distribution preferences.
Appendix F Results of Jailbreaking LLMs
We conduct jailbreaking experiments on Llama 3.1 8B Instruct, Gemma 2B It, Qwen 2 7B Instruct, and Llama 2 7B Chat. The results of preference decomposition, refusal rate decomposition, and comparisons with the original censored models are presented in Figure 12 to Figure 15. Our findings indicate that jailbreaking is more effective on Llama 3.1 8B compared to the other models.
(a) Preference decomposition.
(b) Refusal rate decomposition.
Figure 12: Radar plots depicting the preference decomposition (left) and refusal rate decomposition (right) for Llama 3.1 8B Instruct and its uncensored variant.
(a) Preference decomposition.
(b) Refusal rate decomposition.
Figure 13: Radar plots depicting the preference decomposition (left) and refusal rate decomposition (right) for Gemma 2 2B It and its uncensored variant.
(a) Preference decomposition.
(b) Refusal rate decomposition.
Figure 14: Radar plots depicting the preference decomposition (left) and refusal rate decomposition (right) for Qwen 2 7B Instruct and its uncensored variant.
(a) Preference decomposition.
(b) Refusal rate decomposition.
Figure 15: Radar plots depicting the preference decomposition (left) and refusal rate decomposition (right) for Llama 2 7B Chat and its uncensored variant.